OpenStudy (anonymous):

Find the height of a triangle whose base is 5 inches and area is 12 square inches.

OpenStudy (nali):

Area of a triangle is A=(1/2) b h 12=(1/2) (5) h 12= 2.5 h and just solve for h which is the height